A man is 13 times as old as his son. In ten years he will be 3 times as old as his son will be then. How old is the son now?
Answer by amarjeeth123(570)

Let the son's age be x.
Then the father's age is 13x.
In ten years their ages will be (x+10) and (13x+10).
(13x+10)=3(x+10)
13x+10=3x+30
10x=20
x=2
The son is 2 years old.